OpenStudy (anonymous):

How to express (sin(pi*z^2) + cos (pi*z^2))/((z-1)(z-2)) as partial fractions with z a complex number

OpenStudy (anonymous):

\[\huge\frac{\sin(\pi z^2)+\cos(\pi z^2)}{(z-1)(z-2)}=\frac{\frac{\sin(\pi z^2)}{z-2}}{z-1}+\frac{\frac{\cos(\pi z^2)}{z-1}}{z-2}\] That should work.